Please note the data for a Burrito shop which is counting the number of customers daily.
Brad's Burritos
Period t 1 2 3 4 5 6 7 8 9 10
Observed 108 130 96 144 136 113 120 119 120 106

a. Compute a forecast for the next day by using a 3-day moving average. What is the
forecast? What is the MAD for the last five observations?

b. Compute a forecast for the number of customers for the next day using
exponential smoothing with α = 0.3. What is the forecast? What is the MAD for the last five observations?
Answer #1

The Solution:

(a)

Calculate the forecast using 3-month moving average.

Period

Data

Moving average

1

108

-

2

130

-

3

96

-

4

144

(108 + 130 + 96)/3 = 111.3333

5

136

(130 + 96 + 144)/3 = 123.3333

6

113

(96 + 144 + 136)/3 = 125.3333

7

120

(144 + 136 + 113)/3 = 131

8

119

(136 + 113 + 120)/3 = 123

9

120

(113 + 120 + 119)/3 = 117.3333

10

106

(120 + 119 + 120)/3 = 119.6667

11

-

(119 + 120 + 106)/3 = 115

Now Find Absolute Error for each forecast:

Period

Data

Forecast

Absolute Error

1

108

2

130

3

96

4

144

111.3333

|144- 111.333| = 32.6667

5

136

123.3333

12.6667

6

113

125.3333

12.3333

7

120

131

11

8

119

123

4

9

120

117.3333

2.6667

10

106

119.6667

13.6667

11

-

115

Forecast = 115.00

MAD for last five observations = (12.3333 + 11 + 4 + 2.6667 + 13.6667) / 5 = 43.6667/5 = 8.73334

===============================================================

(b)

Forecast using smoothening constant 0.3

Period

Data

Forecasts

1

108

108

2

130

108+0.3⋅ (108−108) =108

3

96

108+0.3⋅ (130−108) =114.6

4

144

114.6+0.3⋅ (96−114.6) =109.02

5

136

109.02+0.3⋅ (144−109.02) =119.514

6

113

119.514+0.3⋅ (136−119.514) =124.4598

7

120

124.4598+0.3⋅ (113−124.4598) =121.0219

8

119

121.02186+0.3⋅ (120−121.02186) =120.7153

9

120

120.715302+0.3⋅ (119−120.715302) =120.2007

10

106

114+0.3⋅ (120−120.2007114) =120.1405

11

-

120.14049798+0.3⋅ (106−120.14049798) =115.8983

Calculate Absolute Error:

Period

Data

Forecast

Absolute Error

1

108

108

|108-108| = 0

2

130

108

|130 – 108| = 22

3

96

114.6

18.6

4

144

109.02

34.98

5

136

119.514

16.486

6

113

124.4598

11.4598

7

120

121.0219

1.0219

8

119

120.7153

1.7153

9

120

120.2007

0.2007

10

106

120.1405

14.1405

11

-

115.8983

Forecast: 115.8983

MAD for the last five observations = (11.4598 + 1.0219 + 1.7153 + 0.2007 + 14.1405) / 5 = (28.5382)/ 5 = 5.70764